OpenStudy (anonymous):

A bowl of soup at temperature x is left out in a cold room. For every minute that the soup is in the room, its temperature drops by 10t degrees, where t is the number of minutes that the soup has been in the room. If x=150, how many minutes will it take for the soup's temperature to reach 40 degrees?

OpenStudy (anonymous):

40 = 150-10t t = 11 answer: 11 minutes

OpenStudy (anonymous):

\(\ \huge x-10t :40=150-10t : -110=-10t\) \(\ \huge so: 11 min. \).